OpenStudy (anonymous):

If f(x)= 2x 2 + 3, then f(x - 1)= 2x 2 + 4x + 5. A.True B.False

OpenStudy (matt101):

f(x-1)=2(x-1)^2+3 = 2(x^2-2x+1)+3 =2x^2-4x=5 False

OpenStudy (anonymous):

a

OpenStudy (anonymous):

False